# KAN or MLP? Point Cloud Shows the Way Forward

### v2

@@ 3.3 Efficient-KAN
In comparison to KAN, Efficient-KAN utilize rational functions as the base functions within KAN, replacing the B-spline functions. Specifically, each activation function $\phi(x)$ is computed using rational polynomials $P(x)$ and $Q(x)$ as defined in eq. 8, where the degrees of $P(x)$ and $Q(x)$ are denoted as $m$ and $n$, respectively.
@@
where $\{a_{i}\}_{i=1,\cdots,m}$ and $\{b_{j}\}_{j=1,\cdots,n}$ are coefficients of the rational functions, and $w$ is the scaling factor. These parameters are trained via backpropagation, where rational functions implemented with CUDA demonstrate lower computational complexity than B-spline functions and enhanced parallelism. We specifically implement two corresponding optimizations. First, we employed Horner’s method (Horner 1815) for polynomial evaluation to further reduce computational overhead. Horner’s method is described as follows:
@@
$\quad\frac{\partial F}{\partial a_{m}}=\frac{x^{m}}{Q(x)},\quad\frac{\partial F}{\partial b_{n}}=-x^{n}\frac{G(x)}{Q^{3}(x)}P(x),$ (10)
$\quad\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}$
@@
Another significant change in Efficient-KAN is the grouping of input channels, which allows parameter sharing within groups to reduce the number of parameters and computational load. As illustrated in fig. 3, MLP utilizes fixed activation functions, while KAN assigns an activation function to each input-output pair. In contrast, Efficient-KAN integrates the approaches of the former two by sharing activation functions within groups to minimize the number of parameters.  Table 1 presents a quantitative comparison of the parameter counts among the single-layer MLP, the Vanilla KAN, and Efficient-KAN, in which $d_{in}$ and $d_{out}$ represent the input and output dimensions, respectively; $k$ denotes the spline order; $G$ indicates the grid size; $m$ and $n$ correspond to the polynomial degrees of rational functions; and $g$ is the number of groups. The parameter count of Efficient-KAN introduces only a constant term compared to MLP, whereas original KAN’s parameter scale grows roughly as (G+k+2) times that of MLP.
@@ 4 Experiments
In this section, we first introduce detailed settings of experiments. Subsequently, PointKAN is compared and evaluated against various existing methods, especially PointMLP, on multiple downstream tasks, including object classification, part segmentation, and few-shot learning. Finally, we present results of ablation studies.
@@ 4.1 Implementation Details
Experiments are conducted on three benchmarks using a single NVIDIA RTX 4090: ModelNet40 (Wu et al. 2015) and ScanObjectNN (Uy et al. 2019) for classification, and ShapeNetPart (Yi et al. 2016) for part segmentation, with hyperparameters aligned to PointMLP (Ma et al. 2022) for fair comparison. The PointKAN architecture employs four identical stages per task, each containing a GAM, LFP, and GFP module, using a batch size of 18. For classification, SGD optimizes inputs of 1,024 points grouped into 24 per partition; for segmentation, Adam with cosine decay handles 2,048 points (including normals) grouped into 32 per partition. The PointKAN-elite variant increases the batch size to 32 while halving learning rates, retaining other configurations. All ablation studies default to PointKAN and are evaluated primarily on ModelNet40.
Table 2: Classification results on ModelNet40 and ScanObjectNN. The inference model parameters Param.(M), FLOPs(G), and overall accuracy (%) are reported. We compare the KAN-based model with methods based on CNN, GCN, Transformer, Mamba, and MLP.
Method Param.(M) $\downarrow$ FLOPs(G) $\downarrow$ ModelNet40 ScanObjectNN
1k P $\uparrow$ OBJ_BG $\uparrow$ OBJ_ONLY $\uparrow$ PB_T50_RS $\uparrow$
@@
DGCNN (Wang et al. 2019) 1.8 2.4 92.9 82.8 86.2 78.1
PointCNN (Li et al. 2018) 0.6 - 92.2 86.1 85.5 78.5
GBNet (Qiu, Anwar, and Barnes 2021b) 8.8 - 93.8 - - 81.0
MVTN (Hamdi, Giancola, and Ghanem 2021) 11.2 43.7 93.8 92.6 92.3 82.8
@@
Transformer (Vaswani et al. 2017) 22.1 4.8 91.4 79.86 80.55 77.24
PCT (Guo et al. 2021) 2.9 2.3 93.2 - - -
PointMamba (Liang et al. 2024) 12.3 3.6 - 88.30 87.78 82.48
PCM (Zhang et al. 2024) 34.2 45.0 93.4$\pm$0.2 - - 88.1$\pm$0.3
@@
PointNet (Qi et al. 2017a) 3.5 0.5 89.2 73.3 79.2 68.0
PointNet++ (Qi et al. 2017b) 1.5 1.7 90.7 82.3 84.3 77.9
DRNet (Qiu, Anwar, and Barnes 2021a) - - 93.1 - - 80.3
SimpleView (Goyal et al. 2021) - - 93.9 - - 80.5$\pm$0.3
PRA-Net (Cheng et al. 2021) 2.3 93.7 - - 81.0
PointMLP (Ma et al. 2022) 12.6 31.4 93.3 - - 85.0$\pm$0.3
@@
PointNet-KAN (Kashefi 2024) - - 90.5 - - 60.2
PointKAN (Ours) 10.0 9.1 93.7 87.1 86.4 83.5
PointKAN-elite (Ours) 3.1 2.3 93.3 84.3 85.3 84.1
@@
DGCNN (Wang et al. 2019) 31.6 $\pm$ 2.8 40.8 $\pm$ 4.6 19.9 $\pm$ 2.1 16.9 $\pm$ 1.5
DGCNN+OcCo (Wang et al. 2021) 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
DGCNN-CrossPoint (Afham et al. 2022) 92.5 $\pm$ 3.0 94.9 $\pm$ 2.1 83.6 $\pm$ 5.3 87.9 $\pm$ 4.2
@@
Transformer (Vaswani et al. 2017) 87.8 $\pm$ 5.2 93.3 $\pm$ 4.3 84.6 $\pm$ 5.5 89.4 $\pm$ 6.3
MAMBA3D  (Han et al. 2024) 92.6 $\pm$ 3.7 96.9 $\pm$ 2.4 88.1 $\pm$ 5.3 93.1 $\pm$ 3.6
Point-BERT (Yu et al. 2022) 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
MaskPoint (Liu, Cai, and Lee 2022) 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
@@
PointNet (Qi et al. 2017a) 52.0 $\pm$ 3.8 57.8 $\pm$ 4.9 46.6 $\pm$ 4.3 35.2 $\pm$ 4.8
PointNet-CrossPoint (Afham et al. 2022) 90.9 $\pm$ 1.9 93.5 $\pm$ 4.4 86.4 $\pm$ 4.7 90.2 $\pm$ 2.2
PointMLP (Ma et al. 2022) 90.5 $\pm$ 4.2 94.2 $\pm$ 4.1 84.1 $\pm$ 5.6 91.5 $\pm$ 5.1
@@
PointKAN(Ours) 95.9 $\pm$ 3.1 97.9 $\pm$ 2.0 90.5 $\pm$ 4.9 94.0 $\pm$ 3.5
PointKAN-elite(Ours) 95.2 $\pm$ 3.1 96.3 $\pm$ 3.0 88.0 $\pm$ 5.1 93.5 $\pm$ 4.0
4.2 Experimental Results on Downstream Tasks
PointKAN (including PointKAN-elite) demonstrates inherent advantages over PointMLP (Ma et al. 2022) in multi-dimensional benchmarking across downstream tasks. Our models are trained from scratch for each specific task, with no implementation of pre-training methodologies or ensemble voting mechanisms in any experimental configuration.
Synthetic Object Classification on ModelNet40
The ModelNet40 (Wu et al. 2015) dataset constitutes a comprehensive and clean collection of 3D CAD models, encompassing 12,311 human-annotated models spanning 40 distinct categories. The dataset is partitioned into 9,843 models for training and 2,468 models for testing. On ModelNet40, PointKAN achieves 93.7% OA (+0.4% vs PointMLP) with 21% fewer parameters (10M vs 12.6M) and 71% lower FLOPs (9.1G vs 31.4G). Furthermore, PointKAN-elite achieves remarkable parameter efficiency (3.1M parameters) and computational efficiency (2.3G FLOPs), while maintaining competitive performance with 90.7% mAcc and 93.3% OA. Results shown in table 2.
Real-world Object Classification on ScanObjectNN
The ScanObjectNN dataset (Uy et al. 2019), derived from 3D scans of real-world indoor scenes, comprises 15,000 point cloud object samples spanning 15 daily object categories (e.g., chairs, tables, monitors) and encompassing 2,902 distinct object instances. On ScanObjectNN, 15 categories), PointKAN achieves 87.1% (OBJ_BG) and 86.4% (OBJ_ONLY) OA in 200 epochs, outperforming scratch-trained baselines. For PB_T50_RS, it attains 83.5% OA (vs PointMLP’s 85.0%), while PointKAN-elite reaches 84.1% with reduced computation. Results shown in table 2.
Few-shot Learning
Following the $n$-way $m$-shot protocol ($n{\in}\{5,10\}$, $m{\in}\{10,20\}$) on ModelNet40, PointKAN achieves superior few-shot accuracy versus CNN and Transformer baselines (see table 3), outperforming MAMBA3D (Han et al. 2024) and Point-BERT (Yu et al. 2022) in most settings. Excluding the 10-way 10-shot case (marginally below MaskPoint (Liu, Cai, and Lee 2022)), it shows consistent OA gains over PointMLP (+5.4%, +3.7%, +6.4%, and +2.5%). Notably, its scratch-trained performance exceeds Point-BERT’s pre-trained results, demonstrating strong generalization.
Part Segmentation on ShapeNetPart
ShapeNetPart (Yi et al. 2016) to predict more fine-grained category labels for each point within the samples. The ShapeNetPart dataset comprises 16,881 3D instances spanning 16 object categories with a total of 50 part labels. We evaluate part segmentation on ShapeNetPart. As shown in  table 4, PointKAN achieves competitive mIoU versus PointMLP with fewer parameters (13.8M vs 16.0M) and FLOPs (3.8G vs 5.8G). PointKAN-elite further improves efficiency (6.2M vs. 16.0M, 1.4G vs. 5.8G). We also provide a comparative visualization between the segmentation ground truths and predicted outcomes in  fig. 4, demonstrating KAN’s potential. Additional part segmentation visualization results are provided in the supplementary materials.
@@ 4.2 Experimental Results on Downstream Tasks
PCNN (Atzmon, Maron, and Lipman 2018) 81.8 85.1 5.4 -
DGCNN (Wang et al. 2019) 82.3 85.2 1.3 12.4
PointCNN (Li et al. 2018) 84.6 86.1 - -
RS-CNN (Liu et al. 2019) 84.0 86.2 - -
SpiderCNN (Xu et al. 2018) 82.4 85.3 - -
@@
PointNet (Qi et al. 2017a) 80.4 83.7 3.6 4.9
PointNet++ (Qi et al. 2017b) 81.9 85.1 1.0 4.9
PointMLP (Ma et al. 2022) 84.6 86.1 16.0 5.8
@@
PointKAN(Ours) 83.9 85.6 13.8 3.8
PointKAN-elite(Ours) 84.3 86.0 6.2 1.4
4.3 Ablation Study
We perform comprehensive ablation studies on the architecture and KAN parameters. All results are obtained from models trained from scratch on ModelNet40, with additional analyses provided in supplementary materials.
Architecture Ablation
Architecture ablation studies ( table 5) demonstrate that integrating S-Pool in GAM elevates overall accuracy by 0.7% through compensating global information loss during local grouping via parallel processing with Group-Norm; conversely, removing LFP reduces OA by 1.0%, validating KAN’s critical role in local feature extraction for unordered point clouds, while omitting the Residual Point Block in GFP decreases accuracy by 0.5%, confirming its necessity for global feature aggregation. The full component integration achieves peak 93.7% OA.
Ablation of Built-in Parameters in KAN
KAN critically determines model performance. Firstly, we analyze two parameters, $k$ (the order of the B-spline function) and $G$ (the number of grids) in  fig. 5. Higher $k$ enhances curve continuity while $G$ controls knot interval density, jointly affecting flexibility. However, excessive values cause overfitting and computational complexity.
In addition, guided by KAN’s generalization principles  (Liu et al. 2024), we balance network scale: insufficient capacity impairs feature learning, while oversized networks harm generalization. Through architectural experiments ( table 6), a three-layer KAN with $k=3$, $G=5$, and hidden dimension $d_{hid}=d_{in}/2$ achieves optimal accuracy.
5 Conclusion
We propose PointKAN, an efficient KAN-based architecture for point cloud analysis. PointKAN outperforms PointMLP across multiple tasks, demonstrating superior local feature extraction. The lightweight PointKAN-elite variant further maintains accuracy while substantially reducing parameters and FLOPs. We anticipate PointKAN will advance KAN applications in point cloud analysis and leverage its unique advantages over MLP.
S-Pool LFP GFP mAcc(%) OA(%)
90.8 93.0
89.9 92.7
90.7 93.2
91.4 93.7

### v4

@@ 3.3 Efficient-KAN
In comparison to KAN, Efficient-KAN utilize rational functions as the base functions within KAN, replacing the B-spline functions. Specifically, each activation function $\phi(x)$ is computed using rational polynomials $P(x)$ and $Q(x)$ as defined in Equation 6, where the degrees of $P(x)$ and $Q(x)$ are denoted as $m$ and $n$, respectively.
@@
where $\{a_{i}\}_{i=1,\cdots,m}$ and $\{b_{j}\}_{j=1,\cdots,n}$ are coefficients of the rational functions, and $w$ is the scaling factor. These parameters are trained via backpropagation, where rational functions implemented with CUDA demonstrate lower computational complexity than B-spline functions and enhanced parallelism. We specifically implement two corresponding optimizations.
First, we employed Horner’s method (Horner, 1815) for polynomial evaluation to further reduce computational overhead. Horner’s method is described as follows:
@@
$\quad\frac{\partial F}{\partial a_{m}}=\frac{x^{m}}{Q(x)},\quad\frac{\partial F}{\partial b_{n}}=-x^{n}\frac{G(x)}{Q^{3}(x)}P(x),\quad\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}$ (8)
@@
Another significant change in Efficient-KAN is the grouping of input channels, which allows parameter sharing within groups to reduce the number of parameters and computational load. As illustrated in Figure 4, MLP utilizes fixed activation functions, while KAN assigns an activation function to each input-output pair. In contrast, Efficient-KAN integrates the approaches of the former two by sharing activation functions within groups to minimize the number of parameters. Table 1 presents a quantitative comparison of the parameter counts among the single-layer MLP, the Vanilla KAN, and Efficient-KAN, in which $d_{in}$ and $d_{out}$ represent the input and output dimensions, respectively; $k$ denotes the spline order; $G$ indicates the grid size; $m$ and $n$ correspond to the polynomial degrees of rational functions; and $g$ is the number of groups. The parameter count of Efficient-KAN introduces only a constant term compared to MLP, whereas original KAN’s parameter scale grows roughly as (G+k+2) times that of MLP.
@@ 4.1 Implementation Details
Experiments are conducted on three benchmarks using NVIDIA RTX 4090: ModelNet40 (Wu et al., 2015) and ScanObjectNN (Uy et al., 2019) for classification, and ShapeNetPart (Yi et al., 2016) for part segmentation, with hyperparameters aligned to PointMLP (Ma et al., 2022) for fair comparison. The PointKAN architecture employs four identical stages per task, each containing a GAM, LFP, and GFP module, using a batch size of 18. For classification, SGD optimizes inputs of 1,024 points grouped into 24 per partition; for segmentation, Adam with cosine decay handles 2,048 points (including normals) grouped into 32 per partition. The PointKAN-elite variant increases the batch size to 32 while halving learning rates, retaining other configurations. All ablation studies default to PointKAN and are evaluated primarily on ModelNet40.
Table 2: Classification results on ModelNet40 and ScanObjectNN. The model parameters Param.(M), FLOPs(G), overall accuracy(%) and Train/Test speed(samples/s) are reported. We compare the KAN-based model with methods based on CNN, GCN, Transformer, Mamba, and MLP.
Method Param.(M) $\downarrow$ FLOPs(G) $\downarrow$ ModelNet40 ScanObjectNN Train speed Test speed
1k P $\uparrow$ OBJ_BG $\uparrow$ OBJ_ONLY $\uparrow$ PB_T50_RS $\uparrow$ (samples/s) $\uparrow$ (samples/s) $\uparrow$
@@
DGCNN (Wang et al., 2019) 1.8 2.4 92.9 82.8 86.2 78.1 - -
PointCNN (Li et al., 2018) 0.6 - 92.5 86.1 85.5 78.5 - -
PointConv (Wu et al., 2019) 18.6 - 92.5 - - - 17.9 10.2
KPConv (Wu et al., 2019) 15.2 - 92.9 - - - 31.0 80.0
GBNet (Qiu et al., 2021b) 8.8 - 93.3 - - 81.0 - -
@@
Transformer (Vaswani et al., 2017) 22.1 4.8 91.4 79.9 80.6 77.2 - -
PCT Guo et al. (2021) 2.9 2.3 93.2 - - - - -
PointMamba (Liang et al., 2024) 12.3 3.6 - 88.3 87.8 82.5 - -
[…]
@@
PointNet (Qi et al., 2017a) 3.5 0.5 89.2 73.3 79.2 68.0 - -
PointNet++ (Qi et al., 2017b) 1.5 1.7 90.7 82.3 84.3 77.9 223.8 308.5
DRNet (Qiu et al., 2021a) - - 93.1 - - 80.3 - -
[…]
PRA-Net (Cheng et al., 2021) 2.3 93.7 - - 81.0 - -
[…]
@@
PointNet-KAN (Kashefi, 2024) - - 90.5 - - 60.2 - -
[…]
4.2 Experimental Results on Downstream Tasks
PointKAN (including PointKAN-elite) demonstrates inherent advantages over PointMLP (Ma et al., 2022) in multi-dimensional benchmarking across downstream tasks. Our models are trained from scratch for each specific task, with no implementation of pre-training methodologies or ensemble voting mechanisms in any experimental configuration.
Synthetic Object Classification on ModelNet40. The ModelNet40 (Wu et al., 2015) dataset constitutes a comprehensive and clean collection of 3D CAD models, encompassing 12,311 human-annotated models spanning 40 distinct categories. The dataset is partitioned into 9,843 models for training and 2,468 models for testing. As demonstrated in Table 2, we report overall accuracy (OA) on the test set. Among these methods, our PointKAN achieves a 0.4% enhancement in OA (93.7% vs. 93.3%) compared to PointMLP (Ma et al., 2022) when using 1k points. Notably, PointKAN reduces parameter count by 21% (10M vs. 12.6M) and computational complexity (FLOPs) by 71% (9.1G vs. 31.4G) relative to PointMLP. Moreover, PointKAN-elite further reduces the number of parameters (3.1M) and computational cost (2.3G FLOPs) while maintaining accuracy. Additionally, its inference speed surpasses that of PointMLP (154.3 samples/s vs. 112 samples/s).
@@ 4.2 Experimental Results on Downstream Tasks
DGCNN (Wang et al., 2019) 31.6 $\pm$ 2.8 40.8 $\pm$ 4.6 19.9 $\pm$ 2.1 16.9 $\pm$ 1.5
DGCNN+OcCo (Wang et al., 2021) 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
DGCNN-CrossPoint (Afham et al., 2022) 92.5 $\pm$ 3.0 94.9 $\pm$ 2.1 83.6 $\pm$ 5.3 87.9 $\pm$ 4.2
@@
Transformer (Vaswani et al., 2017) 87.8 $\pm$ 5.2 93.3 $\pm$ 4.3 84.6 $\pm$ 5.5 89.4 $\pm$ 6.3
MAMBA3D  (Han et al., 2024) 92.6 $\pm$ 3.7 96.9 $\pm$ 2.4 88.1 $\pm$ 5.3 93.1 $\pm$ 3.6
Point-BERT (Yu et al., 2022) 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
MaskPoint (Liu et al., 2022) 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
@@
PointNet (Qi et al., 2017a) 52.0 $\pm$ 3.8 57.8 $\pm$ 4.9 46.6 $\pm$ 4.3 35.2 $\pm$ 4.8
PointNet-CrossPoint (Afham et al., 2022) 90.9 $\pm$ 1.9 93.5 $\pm$ 4.4 86.4 $\pm$ 4.7 90.2 $\pm$ 2.2
PointMLP (Ma et al., 2022) 90.5 $\pm$ 4.2 94.2 $\pm$ 4.1 84.1 $\pm$ 5.6 91.5 $\pm$ 5.1
@@
[…]
Real-world Object Classification on ScanObjectNN. The ScanObjectNN dataset (Uy et al., 2019), derived from 3D scans of real-world indoor scenes, comprises 15,000 point cloud object samples spanning 15 daily object categories (e.g., chairs, tables, monitors) and encompassing 2,902 distinct object instances. As shown in Table 2, we conducted experiments on three progressively challenging variants of ScanObjectNN: OBJ-BG (objects with background), OBJ-ONLY (isolated objects), and PB-T50-RS (perturbed objects with real-world noise). Among models trained from scratch, PointKAN achieved exceptional overall accuracy (OA) with fewer training epochs (200 epochs), attaining 87.1% on OBJ-BG and 86.4% on OBJ-ONLY. However, on the most complex PB-T50-RS variant, PointKAN slightly underperformed PointMLP (83.5% vs. 84.7% OA). Notably, PointKAN-elite demonstrated a 0.6% OA improvement over PointKAN (84.1% vs. 83.5%) on PB-T50-RS while achieving a significant reduction in parameters and FLOPs.
Few-shot Learning. To validate the few-shot transfer capability of PointKAN, we conduct few-shot classification experiments following prior work. Adopting the ”n-way, m-shot” protocol on ModelNet40, where $n\in\{5,10\}$ denotes the number of randomly selected object categories and $m\in\{10,20\}$ represents the number of sampled instances per category, we restrict model training to $n\times m$ samples exclusively. During testing, 20 novel instances per class are randomly selected as evaluation data. For each configuration, we perform 10 independent trials and report the mean accuracy with standard deviation. As shown in Table 3, PointKAN achieves superior accuracy compared to CNN-based variants and even surpasses Transformer/Mamba-based counterparts like Point-BERT (Yu et al., 2022) and MAMBA3D (Han et al., 2024). The only exception occurs in the 10-way 10-shot setting, where PointKAN marginally underperforms MaskPoint (Liu et al., 2022) and Point-BERT. In comparisons between KANs and MLPs, PointKAN demonstrates consistent overall accuracy (OA) gains over PointMLP: +5.4%, +3.7%, +6.4%, and +2.5% across configurations. Notably, despite being trained from scratch, PointKAN (and its lightweight variant PointKAN-elite) outperforms Point-BERT (which employs self-supervised pre-training) in most tasks, underscoring its exceptional generalization capacity and knowledge transfer ability.
@@
PCNN (Atzmon et al., 2018) 81.8 85.1 5.4 -
DGCNN (Wang et al., 2019) 82.3 85.2 1.3 12.4
PointCNN (Li et al., 2018) 84.6 86.1 - -
RS-CNN (Liu et al., 2019) 84.0 86.2 - -
SpiderCNN (Xu et al., 2018) 82.4 85.3 - -
@@
PointNet (Qi et al., 2017a) 80.4 83.7 3.6 4.9
PointNet++ (Qi et al., 2017b) 81.9 85.1 1.0 4.9
PointMLP (Ma et al., 2022) 84.6 86.1 16.0 5.8
@@
[…]
Part Segmentation on ShapeNetPart. We conduct part segmentation experiments on ShapeNetPart (Yi et al., 2016) to predict more fine-grained category labels for each point within the samples. The ShapeNetPart dataset comprises 16,881 3D instances spanning 16 object categories with a total of 50 part labels. As illustrated in Table 4, we systematically compared our method against representative approaches based on CNNs, GCNs, and MLPs, reporting both the mean Intersection over Union (mIoU) across all categories (Cls.) and instances (Inst.). Due to space limitations, we report the mIoU for each category in the supplementary materials. The experimental results demonstrate that our PointKAN (including PointKAN-elite variant) achieves highly competitive performance compared to PointMLP while exhibiting significant reductions in both parameter count (13.8M vs. 16.0M) and FLOPs(3.8G vs. 5.8G). PointKAN-elite demonstrates even more significant improvements(6.2M vs. 16.0M, 1.4G vs. 5.8G). (Visualization of part segmentation in the appendix.)
Table 5: Impact of components on model robustness. We conduct ablation studies on ModelNet40 and the overall accuracy (%) are reported.
Affine S-Pool Translation and scaling Point dropout Noise
92.5 91.0 90.3
93.0 91.6 90.5
90.7 90.2 90.0
93.7 92.9 91.8